OpenStudy (solomonzelman):

Don't post links please. I need to some correction (if I am wrong). Look at my statements about translating functions and tel me if I am correct or not. Can you also please explain why is each of them true. (I kind of get that, but not for sure.)

OpenStudy (solomonzelman):

Lets say you have a function f(x)=x+3 1) f(x)=(x-a)+3 translated a units to the right 2) f(x)=(x+a)+3 translated a units to the left 3) f(x)=x+3-a translated a units down 4) f(x)=x+3+a translated a units up Are these statements right?

OpenStudy (solomonzelman):

So it is correct?

OpenStudy (solomonzelman):

(apparently yes, since I got a medal)